OpenStudy (anonymous):

f(x) = xsquared -2x + 3 f(x + 4) =?????????

OpenStudy (asnaseer):

just replace all occurrences of x with x+4 in the expression.

OpenStudy (anonymous):

i know but i keep getting the wrong answer

OpenStudy (asnaseer):

show me your steps please and I will try and see where you made a mistake.

OpenStudy (anonymous):

= (x + 4) squared - 2(x + 4) +3 = xsquared + 16 -2x - 8 +3 = xsquared -2x + 11

OpenStudy (asnaseer):

ok - you have expanded \((x+4)^2\) incorrectly. it should be as follows:\[(x+4)^2=(x+4)(x+4)=x(x+4)+4(x+4)=x^2+4x+4x+16\]\[\qquad=x^2+8x+16\]

OpenStudy (anonymous):

k

OpenStudy (asnaseer):

do you get the right answer with this expansion now?